Igor invested $8000 in two accounts paying 1% and 5% annual interest respectively. If total interest earned for the year was $320, how much was invested at each rate?
Found 2 solutions by stanbon, greenestamps:
Answer by stanbon(75887)

Igor invested $8000 in two accounts paying 1% and 5% annual interest respectively. If total interest earned for the year was $320, how much was invested at each rate?
-------------
Equation:
int + int = 320
0.01x + 0.05(8000-x) = 320
----------------------
x + 5*8000-5x = 32000
---------------------------
-4x = -8000
x = 2000 (amt. invested at 1%)
8000-x = 6000 (amt. invested at 5%)
--------
Cheers,
Stan H.
---------

Answer by greenestamps(13215)



Here is an alternative solution method that can get you to the answer faster, if you understand it....

$320 interest on an investment of $8000 is 4%.
The interest rates on the two accounts are 1% and 5%.
4% is "three times as close" to 5% as it is to 1%. (5-4=1; 4-1=3)
Therefore the amount invested at 5% must be 3 times the amount invested at 1%.
So the $8000 must be divided into $6000 at 5% and $2000 at 1%.